# Fractionally modulated discrete Carleson’s Theorem and pointwise Ergodic Theorems along certain curves

### v1

@@ 4. Pointwise Ergodic Theorem
$\big{\|}\sup_{t\in\mathbb{N}}|A_{t}f|\big{\|}_{\ell^{p}(\mathbb{Z}^{2})}\leq C\|f\|_{\ell^{p}(\mathbb{Z}^{2})}\quad\text{and}\quad\|V^{r}(A_{\lfloor\lambda^{n}\rfloor}f:\,n\in\mathbb{N})\|_{\ell^{2}(\mathbb{Z})}\leq C^{\prime}\|f\|_{\ell^{2}(\mathbb{Z})}\text{.}$
@@
We begin with the maximal estimate. Firstly, note that $\sup_{t\in\mathbb{N}}|A_{t}f|\lesssim\sup_{n\in\mathbb{N}_{0}}A_{2^{n}}|f|$. Let $\eta\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{2})$ be such that $\bm{1}_{[-1/4,1/4)^{2}}\leq\eta\leq\bm{1}_{[-1/2,1/2)^{2}}$, and let $\mathcal{L}_{t}(\xi_{1},\xi_{2}):=\frac{1}{t}\int_{0}^{t}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s$. Fix $n_{0}\in\mathbb{N}$ and note that by the monotone convergence theorem it suffices to show that
@@
The first summand is immediately bounded by the previous theorem, and for the second summand, we fix $\varepsilon$ as in the previous section, and for $p=2$, by Proposition 3.6 we have that
$\big{\|}\mathcal{F}^{-1}_{\mathbb{R}^{2}}\big{[}(1-\eta)\mathcal{L}_{2^{n}}\mathcal{F}_{\mathbb{R}^{2}}f\big{]}\big{\|}_{L^{2}(\mathbb{R}^{2})}\leq\|(1-\eta)\mathcal{L}_{2^{n}}\|_{L^{\infty}}\|f\|_{L^{2}(\mathbb{R}^{2})}\lesssim 2^{-\frac{\varepsilon}{4}n}\|f\|_{L^{2}(\mathbb{R}^{2})}\text{,}$
@@
$\big{\|}\mathcal{F}^{-1}_{\mathbb{R}^{2}}\big{[}(1-\eta)\mathcal{L}_{2^{n}}\mathcal{F}_{\mathbb{R}^{2}}f\big{]}\big{\|}_{L^{p}(\mathbb{R}^{2})}\lesssim_{\eta}2^{-\frac{\varepsilon(p,p_{0})}{4}n}\|f\|_{L^{p}(\mathbb{R}^{2})}\text{,}\quad\text{for some $\varepsilon(p,p_{0})>0$.}$
@@
$\Big{\|}\sup_{n\leq n_{0}}\big{|}\mathcal{F}^{-1}_{\mathbb{R}^{2}}\big{[}\eta\mathcal{L}_{2^{n}}\mathcal{F}_{\mathbb{R}^{2}}f\big{]}\big{|}\Big{\|}_{L^{p}(\mathbb{R}^{2})}\lesssim\|f\|_{L^{p}(\mathbb{R}^{2})}+\Big{(}\sum_{n\in\mathbb{N}}2^{-\frac{p\varepsilon(p,p_{0})}{4}n}\|f\|^{p}_{L^{p}(\mathbb{R}^{2})}\Big{)}^{1/p}\lesssim\|f\|_{L^{p}(\mathbb{R}^{2})}\text{.}$
@@
$\big{\|}\mathcal{F}^{-1}_{\mathbb{Z}^{2}}\big{[}(k_{2^{n}}-\eta\mathcal{L}_{2^{n}})\mathcal{F}_{\mathbb{Z}^{2}}f\big{]}\big{\|}_{\ell^{2}(\mathbb{Z}^{2})}\leq\|k_{2^{n}}-\eta\mathcal{L}_{2^{n}}\|_{L^{\infty}([-1/2,1/2))}\|f\|_{\ell^{2}(\mathbb{Z})}\lesssim 2^{-\frac{\varepsilon}{4}n}\|f\|_{\ell^{2}(\mathbb{Z})}\text{,}$
@@
(4.7) $\|k_{2^{n}}-\eta\mathcal{L}_{2^{n}}\|_{L^{\infty}([-1/2,1/2))}\leq\|k_{2^{n}}-\mathcal{L}_{2^{n}}\|_{L^{\infty}(\widetilde{\mathfrak{M}}_{2^{n}})}+\big{\|}|k_{2^{n}}|+|\mathcal{L}_{2^{n}}|\big{\|}_{L^{\infty}([-1/2,1/2)\setminus\widetilde{\mathfrak{M}}_{2^{n}})}\lesssim 2^{-\frac{\varepsilon}{4}n}\text{.}$
@@
$\big{\|}\mathcal{F}^{-1}_{\mathbb{Z}^{2}}\big{[}(k_{2^{n}}-\eta\mathcal{L}_{2^{n}})\mathcal{F}_{\mathbb{Z}^{2}}f\big{]}\big{\|}_{\ell^{p}(\mathbb{Z}^{2})}\lesssim 2^{-\varepsilon(p,p_{0})n}\|f\|_{\ell^{p}(\mathbb{Z}^{2})}\text{,}$
@@
where for the last estimate we used the fact that for large $n$ we have that $\big{\|}k_{\lfloor\lambda^{n}\rfloor}-\mathcal{L}_{\lfloor\lambda^{n}\rfloor}\tilde{\eta}\big{\|}_{L^{\infty}(\mathbb{T}^{2})}\lesssim\lambda^{-\frac{\varepsilon}{4}n}$, and this can be justified in an analogous manner to (4.7). The proof is complete. ∎
@@ 5. Reducing the proof of Theorem 1 to the continuous case
In this section we begin the proof of Theorem 1, essentially reducing it to Theorem 1.1 in [8]. We note that throughout the last two sections the implicit constants may also depend on $p$. We will use the following Sobolev embedding lemma which can be found in [13, Lemma 4.12]; we state it here for the sake of clarity.
@@
Before appealing to this result, we discard all $\lambda$-frequencies in the supremum defining the $c$-modulated Carleson operators; this is the content of the following proposition, the proof of which will be given in the end of the section and is the most technical part of the present work. Similarly to the beginning of Section 3 we fix a parameter $\varepsilon\in\big{(}0,\min\big{\{}\frac{1}{4},2-c\big{\}}\big{)}$ and we additionally fix $\nu<\varepsilon/10$ and define for every $j\in\mathbb{N}$
$X_{j}\coloneqq\{\lambda\in[-1/2,1/2):\,|\lambda|\leq 2^{(-c+\nu)j}\}\text{.}$
@@
Let $\mathsf{i}\in\{0,1\}$. There exist constants $C$ and $\rho>0$ such that for every $j\in\mathbb{N}$ we have
(5.1) $\Big{\|}\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\big{|}\big{(}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\big{|}\Big{\|}_{\ell^{2}_{dx}}\leq C2^{-\rho j}\|f\|_{\ell^{2}}\text{.}$
@@
(5.2) $\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\sum_{n\neq 0}f(x-n)\frac{e(\lambda\mathsf{sign}^{\mathsf{i}}(n)\lfloor|n|^{c}\rfloor)}{n}\Big{|}\bigg{\|}_{{\ell_{dx}^{p}}}\\ \lesssim\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\Big{(}\sum_{j\in\mathbb{N}}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}+\sum_{j\in\mathbb{N}}\bigg{\|}\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\Big{|}\big{(}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}\\ \lesssim\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\Big{(}\sum_{j\in\mathbb{N}}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}+\|f\|_{\ell^{p}}\text{,}$
where to obtain the last inequality we argue as follows. For $p=2$, (5.1) immediately yields the desired estimate. We note that the following pointwise estimate holds
@@
where $Mf$ is the standard discrete Hardy–Littlewood maximal operator. We obtain weak-type $(1,1)$ and strong (,)($\infty$,$\infty$)( ∞ , ∞ ) estimates and interpolation with the $\ell^{2}$-estimates of (5.1) yields
(5.4) $\bigg{\|}\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\Big{|}\big{(}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}\lesssim 2^{-\rho_{p}j}\|f\|_{\ell^{p}}\text{, for some $\rho_{p}>0$,}$
@@
(5.5) $\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\Big{(}\sum_{j\in\mathbb{N}}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}\\ \lesssim\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\Big{(}\sum_{j\in\mathbb{N}}\hat{f}(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}+\sum_{j\in\mathbb{N}}\bigg{\|}\sup_{\lambda\in X_{j}}\big{|}\big{(}\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\big{|}\bigg{\|}_{\ell_{dx}^{p}}\text{,}$
@@
(5.6) $\bigg{\|}\sup_{\lambda\in X_{j}}\big{|}\big{(}\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\big{|}\bigg{\|}_{\ell_{dx}^{2}}\lesssim\big{(}2^{-\frac{\varepsilon}{4}j}+(2^{-\frac{\varepsilon}{4}j}2^{cj}2^{(-c+\nu)j})^{1/2}\big{)}\|f\|_{\ell^{2}}\lesssim 2^{-\frac{\varepsilon}{16}j}\|f\|_{\ell^{2}}\text{,}\quad\text{since $\nu<\varepsilon/10$.}$
For $p=2$ we can immediately use the above estimate to bound the second summand in the last line of (5.5). For $p\neq 2$ we use interpolation between the $\ell^{2}$-estimates and trivial estimates to obtain an estimate of the form
(5.7) $\bigg{\|}\sup_{\lambda\in X_{j}}\big{|}\big{(}\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\big{|}\bigg{\|}_{\ell_{dx}^{p}}\lesssim 2^{-\rho_{p}j}\|f\|_{\ell^{p}}\text{,}\quad\text{for some $\rho_{p}>0$.}$
@@
$\bigg{\|}\sup_{\lambda\in X_{j}}\big{|}\big{(}\hat{f}(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\big{)}^{\vee}(x)\big{|}\bigg{\|}_{\ell_{dx}^{p_{0}}}\lesssim_{p_{0}}\|f\|_{\ell^{p_{0}}}\text{,}$
@@
$\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\Big{(}\sum_{j\in\mathbb{N}}\hat{f}(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)1_{X_{j}}(\lambda)\Big{)}^{\vee}(x)\Big{|}\bigg{\|}_{\ell_{dx}^{p}}\lesssim\|f\|_{\ell^{p}}\text{.}$
@@
(5.8) $\bigg{\|}\sup_{\lambda\in[-1/2,1/2)}\Big{|}\mathcal{F}^{-1}_{\mathbb{R}}\Big{[}\sum_{j\in\mathbb{N}}\mathcal{F}_{\mathbb{R}}[f](\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)1_{X_{j}}(\lambda)\Big{]}\Big{|}\bigg{\|}_{L^{p}(\mathbb{R})}\lesssim\|f\|_{L^{p}(\mathbb{R})}\text{,}$
@@
We end this section by proving Proposition 5.2.
@@
We remove the floor function. For $n\neq 0$, $\lambda\in[-1/2,1/2)$ and $M\in\mathbb{N}$ we have
@@
as in the beginning of the proof of Proposition 2.1. Fix $\delta_{1}\in\big{(}0,\frac{2-c}{4}\big{)}$ and apply the previous estimate with $M=\lfloor 2^{\delta_{1}j}\rfloor$. We get

### v2

@@ 4. Pointwise Ergodic Theorem
$\big\|\sup_{t\in\mathbb{N}}|A_{t}f|\big\|_{\ell^{p}(\mathbb{Z}^{2})}\leq C\|f\|_{\ell^{p}(\mathbb{Z}^{2})}\quad\text{and}\quad\|V^{r}(A_{\lfloor\lambda^{n}\rfloor}f:\,n\in\mathbb{N})\|_{\ell^{2}(\mathbb{Z}^{2})}\leq C^{\prime}\|f\|_{\ell^{2}(\mathbb{Z}^{2})}\text{.}$
@@
Again we fix $c$ and $\tau\in(0,\min\{1/4,2-c\})$ as before and all implied constants may depend on them. We begin with the maximal estimate. Firstly, note that $\sup_{t\in\mathbb{N}}|A_{t}f|\lesssim\sup_{n\in\mathbb{N}_{0}}A_{2^{n}}|f|$. Let $\eta\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{2})$ be such that $\bm{1}_{[-1/4,1/4)^{2}}\leq\eta\leq\bm{1}_{[-1/2,1/2)^{2}}$, and let $\mathcal{L}_{t}(\xi_{1},\xi_{2}):=\frac{1}{t}\int_{0}^{t}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s$. Fix $n_{0}\in\mathbb{N}$ and note that by the monotone convergence theorem it suffices to show that
@@
The first summand is immediately bounded by the previous theorem, and for the second summand, by Proposition 3.6 we have that
$\big\|\mathcal{F}^{-1}_{\mathbb{R}^{2}}\big[(1-\eta)\mathcal{L}_{2^{n}}\mathcal{F}_{\mathbb{R}^{2}}f\big]\big\|_{L^{2}(\mathbb{R}^{2})}\leq\|(1-\eta)\mathcal{L}_{2^{n}}\|_{L^{\infty}}\|f\|_{L^{2}(\mathbb{R}^{2})}\lesssim 2^{-\frac{\tau}{4}n}\|f\|_{L^{2}(\mathbb{R}^{2})}\text{,}$
@@
$\big\|\mathcal{F}^{-1}_{\mathbb{R}^{2}}\big[(1-\eta)\mathcal{L}_{2^{n}}\mathcal{F}_{\mathbb{R}^{2}}f\big]\big\|_{L^{p}(\mathbb{R}^{2})}\lesssim_{\eta}2^{-\frac{\tau(p,p_{0})}{4}n}\|f\|_{L^{p}(\mathbb{R}^{2})}\text{,}\quad\text{for some $\tau(p,p_{0})>0$.}$
@@
$\Big\|\sup_{n\leq n_{0}}\big|\mathcal{F}^{-1}_{\mathbb{R}^{2}}\big[\eta\mathcal{L}_{2^{n}}\mathcal{F}_{\mathbb{R}^{2}}f\big]\big|\Big\|_{L^{p}(\mathbb{R}^{2})}\lesssim\|f\|_{L^{p}(\mathbb{R}^{2})}+\Big(\sum_{n\in\mathbb{N}}2^{-\frac{p\tau(p,p_{0})}{4}n}\|f\|^{p}_{L^{p}(\mathbb{R}^{2})}\Big)^{1/p}\lesssim\|f\|_{L^{p}(\mathbb{R}^{2})}\text{.}$
@@
$\big\|\mathcal{F}^{-1}_{\mathbb{Z}^{2}}\big[(k_{2^{n}}-\eta\mathcal{L}_{2^{n}})\mathcal{F}_{\mathbb{Z}^{2}}f\big]\big\|_{\ell^{2}(\mathbb{Z}^{2})}\leq\|k_{2^{n}}-\eta\mathcal{L}_{2^{n}}\|_{L^{\infty}([-1/2,1/2))}\|f\|_{\ell^{2}(\mathbb{Z})}\lesssim 2^{-\frac{\tau}{4}n}\|f\|_{\ell^{2}(\mathbb{Z})}\text{,}$
@@
(4.7) $\|k_{2^{n}}-\eta\mathcal{L}_{2^{n}}\|_{L^{\infty}([-1/2,1/2))}\leq\|k_{2^{n}}-\mathcal{L}_{2^{n}}\|_{L^{\infty}(\widetilde{\mathfrak{M}}_{2^{n}})}+\big\||k_{2^{n}}|+|\mathcal{L}_{2^{n}}|\big\|_{L^{\infty}([-1/2,1/2)\setminus\widetilde{\mathfrak{M}}_{2^{n}})}\lesssim 2^{-\frac{\tau}{4}n}\text{.}$
@@
$\big\|\mathcal{F}^{-1}_{\mathbb{Z}^{2}}\big[(k_{2^{n}}-\eta\mathcal{L}_{2^{n}})\mathcal{F}_{\mathbb{Z}^{2}}f\big]\big\|_{\ell^{p}(\mathbb{Z}^{2})}\lesssim 2^{-\tau(p,p_{0})n}\|f\|_{\ell^{p}(\mathbb{Z}^{2})}\text{,}$
@@
where for the last estimate we used the fact that for large $n$ we have that $\big\|k_{\lfloor\lambda^{n}\rfloor}-\mathcal{L}_{\lfloor\lambda^{n}\rfloor}\tilde{\eta}\big\|_{L^{\infty}(\mathbb{T}^{2})}\lesssim\lambda^{-\frac{\tau}{4}n}$, and this can be justified in an analogous manner to (4.7). The proof is complete. ∎
@@ 5. Reducing the proof of Theorem 1 to the continuous case
In this section we begin the proof of Theorem 1, essentially reducing it to Theorem 1.1 in [8]. We will use the following Sobolev embedding lemma which can be found in [13, Lemma 4.12]; we state it here for the sake of clarity.
@@
Before appealing to this result, we discard all $\lambda$-frequencies in the supremum defining the $c$-modulated Carleson operators; this is the content of the following proposition, the proof of which will be given in the end of the section and is the most technical part of the present work. Similarly to the beginning of Section 3 we fix for convenience two parameters $\varepsilon\coloneqq\frac{2-c}{4}\in\big(0,\min\big\{\frac{1}{4},2-c\big\}\big)$ and $\nu\coloneqq\frac{\varepsilon}{8}$ and define for every $j\in\mathbb{N}$
$X_{j}\coloneqq\big\{\lambda\in[-1/2,1/2):\,|\lambda|\leq 2^{(-c+\nu)j}\big\}\text{.}$
@@
There exist an absolute positive constant $C$ such that for every $j\in\mathbb{N}$ and $\mathsf{i}\in\{0,1\}$ we have
(5.1) $\Big\|\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\big|\big(\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\Big\|_{\ell^{2}_{{\rm d}x}}\leq C(c-1)^{-\frac{1}{2}}(2-c)^{-\frac{9}{4}}2^{-\frac{\varepsilon}{256}j}\|f\|_{\ell^{2}}\text{.}$
@@
(5.2) $\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\sum_{n\neq 0}f(x-n)\frac{e(\lambda\mathsf{sign}^{\mathsf{i}}(n)\lfloor|n|^{c}\rfloor)}{n}\Big|\bigg\|_{{\ell_{{\rm d}x}^{p}}}\\ \leq\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\Big(\sum_{j\in\mathbb{N}}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}+\sum_{j\in\mathbb{N}}\bigg\|\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\Big|\big(\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}\\ \lesssim_{c,p}\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\Big(\sum_{j\in\mathbb{N}}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}+\|f\|_{\ell^{p}}\text{,}$
where to obtain the last inequality we argue as follows. For $p=2$, (5.1) yields that
(5.3) $\sum_{j\in\mathbb{N}}\Big\|\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\big|\big(\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\Big\|_{\ell^{2}_{{\rm d}x}}\\ \lesssim(c-1)^{-\frac{1}{2}}(2-c)^{-\frac{9}{4}}\Big(\sum_{j\in\mathbb{N}}2^{-\frac{2-c}{256}j}\Big)\|f\|_{\ell^{2}}\lesssim(c-1)^{-\frac{1}{2}}(2-c)^{-\frac{13}{4}}\|f\|_{\ell^{2}}\text{,}$
where the implicit constant is absolute and we used that for all $\theta\in(0,1)$ we have
(5.4) $\sum_{j\in\mathbb{N}}2^{-\theta j}\leq\frac{1}{1-2^{-\theta}}=\frac{1}{1-e^{-\theta\log(2)}}\leq\frac{e}{\theta\log(2)}\lesssim\theta^{-1}\text{, since $1-e^{-x}\geq\frac{x}{e}$ for $x\in[0,1]$. }$
We note that the following pointwise estimate holds
@@
where $Mf$ is the standard discrete Hardy–Littlewood maximal operator. We obtain weak-type $(1,1)$ and strong $(\infty,\infty)$ estimates and interpolation with the $\ell^{2}$-estimates of (5.1) yields
(5.6) $\bigg\|\sup_{\lambda\in[-1/2,1/2)\setminus X_{j}}\Big|\big(\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}\lesssim_{c,p}2^{-\rho_{p}j}\|f\|_{\ell^{p}}\text{, for some $\rho_{p}>0$,}$
@@
(5.7) $\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\Big(\sum_{j\in\mathbb{N}}\hat{f}(\cdot)m_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}\\ \leq\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\Big(\sum_{j\in\mathbb{N}}\hat{f}(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}+\sum_{j\in\mathbb{N}}\bigg\|\sup_{\lambda\in X_{j}}\big|\big(\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\bigg\|_{\ell_{{\rm d}x}^{p}}\text{,}$
@@
(5.8) $\frac{\bigg\|\sup_{\lambda\in X_{j}}\big|\big(\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\bigg\|_{\ell_{dx}^{2}}}{(c-1)^{-\frac{1}{2}}}\lesssim\big(2^{-\frac{\varepsilon}{4}j}+(2^{-\frac{\varepsilon}{4}j}2^{cj}2^{(-c+\nu)j})^{1/2}\big)\|f\|_{\ell^{2}}\lesssim 2^{-\frac{\varepsilon}{16}j}\|f\|_{\ell^{2}}\text{,}$
since $\nu=\varepsilon/8$. For $p=2$ we can immediately use the above estimate to bound the second summand in the last line of (5.7) as follows
(5.9) $\sum_{j\in\mathbb{N}}\bigg\|\sup_{\lambda\in X_{j}}\big|\big(\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\bigg\|_{\ell_{dx}^{2}}\lesssim(c-1)^{-\frac{1}{2}}\Big(\sum_{j=0}^{\infty}2^{-\frac{\varepsilon}{16}j}\Big)\|f\|_{\ell^{2}}\lesssim(2-c)^{-1}(c-1)^{-\frac{1}{2}}\|f\|_{\ell^{2}}\text{,}$
where we took into account that $\varepsilon=(2-c)/4$ and used a calculation as in (5.4).
For $p\neq 2$ we use interpolation between the $\ell^{2}$-estimates and trivial estimates to obtain an estimate of the form
(5.10) $\bigg\|\sup_{\lambda\in X_{j}}\big|\big(\hat{f}(\cdot)E_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\bigg\|_{\ell_{dx}^{p}}\lesssim_{c,p}2^{-\rho_{p}j}\|f\|_{\ell^{p}}\text{, for some $\rho_{p}>0$,}$
@@
$\bigg\|\sup_{\lambda\in X_{j}}\big|\big(\hat{f}(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\big)^{\vee}(x)\big|\bigg\|_{\ell_{dx}^{p_{0}}}\lesssim_{c,p_{0}}\|f\|_{\ell^{p_{0}}}\text{,}$
@@
$\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\Big(\sum_{j\in\mathbb{N}}\hat{f}(\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)1_{X_{j}}(\lambda)\Big)^{\vee}(x)\Big|\bigg\|_{\ell_{{\rm d}x}^{p}}\lesssim_{c,p}\|f\|_{\ell^{p}}\text{.}$
@@
(5.11) $\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\mathcal{F}^{-1}_{\mathbb{R}}\Big[\sum_{j\in\mathbb{N}}\mathcal{F}_{\mathbb{R}}[f](\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)1_{X_{j}}(\lambda)\Big]\Big|\bigg\|_{L^{p}(\mathbb{R})}\lesssim_{c,p}\|f\|_{L^{p}(\mathbb{R})}\text{,}$
@@
[…]
(5.12) $\sup_{f\in\ell^{2}:\,\|f\|_{\ell^{2}}\leq 1}\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\sum_{n\neq 0}f(x-n)\frac{e(\lambda\mathsf{sign}^{\mathsf{i}}(n)\lfloor|n|^{c}\rfloor)}{n}\Big|\bigg\|_{{\ell_{{\rm d}x}^{2}}}\lesssim(c-1)^{-\frac{1}{2}}(2-c)^{-\frac{13}{4}}\\ +\sup_{f\in L^{2}(\mathbb{R}):\,\|f\|_{L^{2}(\mathbb{R})}\leq 1}\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\mathcal{F}^{-1}_{\mathbb{R}}\Big[\sum_{j\in\mathbb{N}}\mathcal{F}_{\mathbb{R}}[f](\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)1_{X_{j}}(\lambda)\Big]\Big|\bigg\|_{L^{2}(\mathbb{R})}\text{,}$
and unsuprisingly the constant from the continuous counterpart corresponding to the second summand above does not introduce any blowup as $c\to 2^{-}$ and is in fact $\lesssim(c-1)^{-\frac{1}{2}}$ as will be apparent from the proof of Proposition 6.1 is Section 6. We end this section by proving Proposition 5.2.
@@
As in the beginning of Lemma 3.3 it suffices to prove the estimate for $j\in\mathbb{N}$ such that $(2-c)j\geq 2^{20}$, since for the rest it is immediate. We remove the floor function. For $n\neq 0$, $\lambda\in[-1/2,1/2)$ and $M\in\mathbb{N}$ we have
@@
as in the beginning of the proof of Proposition 2.1. We apply the previous estimate with $M=\lfloor 2^{\varepsilon j}\rfloor$ and obtain